# B4P: Simultaneous Grasp and Motion Planning for Object Placement

### v1

@@ II-2 Sampling-based Motion Planning
Motion planning computes a collision-free path in the robot’s configuration space to connect the grasp with the placement configuration in pick and place systems. Sampling-based planners [17] have proven efficient and generalizable in exploring the high-dimensional configuration space of manipulators. Significant acceleration can be obtained through parallelism [18, 19, 20, 21, 22] on both GPU and CPU-based computing. Key properties of sampling-based planners include probabilistic completeness [23] and Voronoi bias [24]. The former ensures that a solution will eventually be found as long as it exists in the search space, while the latter contributes to exploration efficiency in high-dimensional spaces. Based on the the advantages of sampling-based approaches, and along with a novel forest-based planning strategy, the proposed B4P can simultaneously pair valid grasp and placement configurations while generating feasible motion plans to connect them.
@@ III-B Trees and Forests
Same as in most sampling-based motion planning approaches, we use trees to represent the exploration of the valid robot configurations and robot motions to connect between them in the configuration space $\mathcal{Q}$. In our bidirectional framework, we denote by $\mathcal{T}^{pick}_{i}$ a motion tree rooted at the $i-$th sampled picking configuration, and by $\mathcal{T}^{place}_{j}$ a motion tree rooted at the $j-$th sampled placement configuration. In the trees, every node is a collision-free robot configuration and every edge is a collision free path. To enable all trees to simultaneously explore the configuration space and find potential valid pairings between $\mathcal{T}^{pick}_{i}$ and $\mathcal{T}^{place}_{j}$, we collect the trees into two motion forests $\mathcal{F}^{pick}=\{\mathcal{T}^{pick}_{i}\}_{i=1:n}$ and $\mathcal{F}^{place}=\{\mathcal{T}^{place}_{j}\}_{j=1:m}$, with $n$ and $m$ trees respectively. While the trees grow from two forests toward each other, only connections between trees from different forests are allowed.
@@ IV-A Algorithm Overview
Figure 4: [2D Maze] The robot needs to find a feasible path to carry the grasped object from the start pose (left bottom corner) to the target region (red area in the right top corner); [Stick on Shelf] The manipulator needs to grasp the stick and place it on the target region on shelf; [Grocery Shelf] The manipulator needs to grasp the target object on the shelf in a grocery scenarios; [Example Picking Configurations] Left: Sampled grasp poses on each edge of the planar object; Middle: Sampled grasp poses on the green stick; Right: Sampled grasp poses on the spam object.
@@ IV-C Path Repair
When $\textsc{ParaLocalReplan}(begin,end)$ finishes, it is possible that there is no path repair solution can be found. In that case, Alg. 4 will return an empty path to let B4P know that the search needs to continue. Otherwise, B4P will return with a successfully found path $\pi$ for the pick and place task (line #7 in Alg. 1). An example path repair procedure is visualized for a 2D mobile robot pick and place task in Fig. 3.
@@ V-A2 Baselines
To provide a comprehensive evaluation, we introduce two bidirectional RRT-based parallel algorithms as our baseline to compare with the proposed B4P approach.
  • RRT-Individual: Each thread performs bidirectional RRT using its own individual sampled grasp.
  • RRT-Shared: All threads concurrently perform bidirectional RRT using a single shared sampled grasp. In this approach, we set a small time limit and when it expires, the planner restarts and uses a new grasp.
@@ V-A3 System Environment
To further demonstrate our approach’s effectiveness, we evaluate it with more realistic and challenging task environments in 3D.
@@ V-C2 Grocery Shelf Task
In this task, the target object, the can of spam, must be placed in the goal region in an upward orientation. Compared with the stick-on-shelf problem, additional obstacles are added to the environment to simulate the challenging narrow passage. Three different grasp pose examples are visualized in the right figure from the bottom row of Fig. 4. Such an intricate task environment creates a narrow path in the configuration space for the pick and place task, only the horizontal grasp pose can satisfy a placement configuration simultaneously with a feasible motion trajectory as displayed in the solution path in Fig. 4. The experimental result is listed in Table. III. As the proposed B4P framework expands unbiased different trees that can fully explore the configuration space, it outperforms both baselines by a large margin and guarantees a feasible trajectory plan in the given time budget. Meanwhile, we also observe a superlinear speed when increasing the threads for computing as shown in Fig. 6. This feature possesses an inherent advantage for accelerating scalable applications in the real world.

### v2

@@ II-2 Sampling-based Motion Planning
Motion planning computes a collision-free path in the robot’s configuration space to connect the grasp with the placement configuration in pick and place systems. Sampling-based planners [17] have proven efficient and generalizable in exploring the high-dimensional configuration space of manipulators. Significant acceleration can be obtained through parallelism [18, 19, 20, 21, 22] on both GPU and CPU-based computing. Key properties of sampling-based planners include probabilistic completeness [23] and Voronoi bias [24]. The former ensures that a solution will eventually be found as long as it exists in the search space, while the latter contributes to exploration efficiency in high-dimensional spaces. Based on the advantages of sampling-based approaches, and along with a novel forest-based planning strategy, the proposed B4P can simultaneously pair valid grasp and placement configurations while generating feasible motion plans to connect them.
@@ III-B Trees and Forests
Same as in most sampling-based motion planning approaches, we use trees to represent the exploration of the valid robot configurations and robot motions to connect between them in the configuration space $\mathcal{Q}$. In our bidirectional framework, we denote by $\mathcal{T}^{pick}_{i}$ a motion tree rooted at the $i-$th sampled picking configuration, and by $\mathcal{T}^{place}_{j}$ a motion tree rooted at the $j-$th sampled placement configuration. We also denote $\mathcal{T}^{pick}_{i}.pick$ to be the $SE(3)$ transformation of the gripper relative to the object. In the trees, every node is a collision-free robot configuration and every edge is a collision free path. To enable all trees to simultaneously explore the configuration space and find potential valid pairings between $\mathcal{T}^{pick}_{i}$ and $\mathcal{T}^{place}_{j}$, we collect the trees into two motion forests $\mathcal{F}^{pick}=\{\mathcal{T}^{pick}_{i}\}_{i=1:n}$ and $\mathcal{F}^{place}=\{\mathcal{T}^{place}_{j}\}_{j=1:m}$, with $n$ and $m$ trees respectively. While the trees grow from two forests toward each other, only connections between trees from different forests are allowed.
@@ IV-A Algorithm Overview
Figure 4: [2D Maze] The robot needs to find a feasible path to carry the grasped object from the start pose (left bottom corner) to the target region (red area in the right top corner); [Stick on Shelf] The manipulator needs to grasp the stick and place it on the target region on shelf; [Grocery Shelf] The manipulator needs to grasp the target object on the shelf in a grocery scenarios; [Example Picking Configurations] Left: Sampled grasp poses on each edge of the planar object; Middle: Sampled grasp poses on the green stick; Right: Sampled grasp poses on the spam object. For each task, grasp poses are uniformly randomly sampled. The optimal number of pick and place trees generated depends on the task.
@@ IV-C Path Repair
When $\textsc{ParaLocalReplan}(begin,end)$ finishes, it is possible that there is no path repair solution can be found. In that case, Alg. 4 will return an empty path to let B4P know that the search needs to continue. Otherwise, B4P will return with a successfully found path $\pi$ for the pick and place task (line #7 in Alg. 1). An example path repair is visualized for a 2D mobile robot pick and place task in Fig. 3.
@@ V-A2 Baselines
To provide a comprehensive evaluation, we introduce two bidirectional RRT parallel algorithms as our baseline to compare with the proposed B4P approach. These planners sample 1 placement with their sampled grasp and only use one placement as a goal.
  • RRT-Individual: Each thread $t$ performs bidirectional RRT using its own individual sampled grasp $q^{pick}_{t}$.
  • RRT-Shared: All threads perform one bidirectional RRT in parallel using a single grasp. In this approach, we set a small time limit and when it expires, the planner restarts and uses a new grasp.
@@ V-A3 System Environment
To further demonstrate our approach’s effectiveness, we evaluate it with more realistic and challenging task environments in 3D. Our results for the speedup of B4P are compared against using B4P with only one thread.
@@ V-C2 Grocery Shelf Task
In this task, the target object, the can of spam, must be placed in the goal region in an upward orientation. Compared with the stick-on-shelf problem, additional obstacles are added to the environment to simulate the challenging narrow passage. Three different grasp pose examples are visualized in the right figure from the bottom row of Fig. 4. Such an intricate task environment creates a narrow path in the configuration space for the pick and place task, only the horizontal grasp pose can satisfy a placement configuration simultaneously with a feasible motion trajectory as displayed in the solution path in Fig. 4. The experimental result is listed in Table. III. As the proposed B4P framework expands unbiased different trees that can fully explore the configuration space, it outperforms both baselines by a large margin and guarantees a feasible trajectory plan in the given time budget. Meanwhile, we also observe a superlinear speed when increasing the threads for computing as shown in Fig. 6. This feature possesses an inherent advantage for accelerating scalable applications in the real world. The speedup observed for this task is lower than that of the Stick-on-Shelf task, likely due to the greater number of valid grasps available for this task.